# An Economic Solution to Copyright Challenges of Generative AI

### v2

@@ Acknowledgements
JTW and ZD conducted this work as independent researchers. HCO, BB, and WJS were supported in part by NSF grant DMS-2310679, Wharton AI for Business, a Simons Investigator Fellowship, the Simons Foundation Math+X Grant to the University of Pennsylvania, NSF grant DMS-2134157, DARPA grant W911NF2010021, and DOE grant DE-SC0022199. This work was done while BB was at Harvard University. We are grateful to Peter Henderson for providing helpful feedback on an early version of this paper.
@@ Appendix C Experimental Settings
The generative models used in our experiment follow from the recent advancements in high-resolution image synthesis utilizing latent diffusion models [28]. [^2] Part of the codebase is adapted from https://github.com/artem-gorodetskii/WikiArt-Latent-Diffusion and https://github.com/VSehwag/minimal-diffusion. All images were cropped and resized to $512\times 512$ resolution. As training a new generative model from scratch would be prohibitively costly, we use LoRA [14], an efficient fine-tuning method that enables us to scale up the models being used in a tractable manner.333 The implementation is adapted from https://huggingface.co/blog/lora. The model used for fine-tuning is Stable Diffusion V1-$4$https://huggingface.co/CompVis/stable-diffusion-v1-4., which is a latent image diffusion model trained on LAION2B-en [^5] https://huggingface.co/datasets/laion/laion2B-en.. For the WikiArt dataset, we train each model on painting images from the same artist, with the text prompt “A painting in the style of [artist name].” Similarly, for the FlickrLogo-27 dataset, we train each model on logo images from the same artist, with the text prompt “A logo by [company name].” For each dataset, the model is fine-tuned with an initial learning rate $10^{-4}$, 10 epochs, and batch size $4$. We train the diffusion models on A100 GPU cluster. For all the SRS results in the maintext, they are being averaged over 20 $x^{(\text{gen})}$’s that are randomly selected/generated according to the specifications in the maintext.
@@
where $\mu_{\theta}(x_{t},t)$ and $\Sigma_{\theta}(x_{t},t)$ are the mean and covariance of the Gaussian distribution at step $t$, learned by the model with parameters $\theta$666 We note that for text-to-image diffusion models, $\mu_{\theta}(\cdot)$ and $\Sigma_{\theta}(\cdot)$ will also depend on the input text encoding, but the core mathematical representation remains the same. The model is trained to minimize the difference between the noisy data and its prediction of the denoised data at each step. This can be formalized as minimizing a loss function, e.g., the mean squared error (MSE), between the original data and the reconstructed data:

### v3

@@ Acknowledgements
JTW and ZD conducted this work as independent researchers. HCO, BB, and WJS were supported in part by NSF grant DMS-2310679, Wharton AI for Business, a Simons Investigator Fellowship, the Simons Foundation Math+X Grant to the University of Pennsylvania, NSF grant DMS-2134157, DARPA grant W911NF2010021, and DOE grant DE-SC0022199. We are grateful to Peter Henderson for providing helpful feedback on an early version of this paper.
@@ Appendix C Experimental Settings
The generative models used in our experiment follow from the recent advancements in high-resolution image synthesis utilizing latent diffusion models [28]. [^3] Part of the codebase is adapted from https://github.com/artem-gorodetskii/WikiArt-Latent-Diffusion and https://github.com/VSehwag/minimal-diffusion. All images were cropped and resized to $512\times 512$ resolution. As training a new generative model from scratch would be prohibitively costly, we use LoRA [14], an efficient fine-tuning method that enables us to scale up the models being used in a tractable manner.444 The implementation is adapted from https://huggingface.co/blog/lora. The model used for fine-tuning is Stable Diffusion V1-4555https://huggingface.co/CompVis/stable-diffusion-v1-4., which is a latent image diffusion model trained on LAION2B-en [^6] https://huggingface.co/datasets/laion/laion2B-en.. For the WikiArt dataset, we train each model on painting images from the same artist, with the text prompt “A painting in the style of [artist name].” Similarly, for the FlickrLogo-27 dataset, we train each model on logo images from the same artist, with the text prompt “A logo by [company name].” For each dataset, the model is fine-tuned with an initial learning rate $10^{-4}$, 10 epochs, and batch size $4$. We train the diffusion models on A100 GPU cluster. For all the SRS results in the maintext, they are being averaged over 20 $x^{(\text{gen})}$’s that are randomly selected/generated according to the specifications in the maintext.
@@
where $\mu_{\theta}(x_{t},t)$ and $\Sigma_{\theta}(x_{t},t)$ are the mean and covariance of the Gaussian distribution at step $t$, learned by the model with parameters $\theta$777 We note that for text-to-image diffusion models, $\mu_{\theta}(\cdot)$ and $\Sigma_{\theta}(\cdot)$ will also depend on the input text encoding, but the core mathematical representation remains the same. The model is trained to minimize the difference between the noisy data and its prediction of the denoised data at each step. This can be formalized as minimizing a loss function, e.g., the mean squared error (MSE), between the original data and the reconstructed data: